Consider the titration of a 25.0 mL sample of 0.180 molL−1 CH3NH2 (Kb=4.4×10−4) with 0.155 molL−1 HBr.

Determine the pH at 5.0 mL of added acid
Answer #1

Given:

M(HBr) = 0.155 M

V(HBr) = 5 mL

M(CH3NH2) = 0.18 M

V(CH3NH2) = 25 mL

mol(HBr) = M(HBr) * V(HBr)

mol(HBr) = 0.155 M * 5 mL = 0.775 mmol

mol(CH3NH2) = M(CH3NH2) * V(CH3NH2)

mol(CH3NH2) = 0.18 M * 25 mL = 4.5 mmol

We have:

mol(HBr) = 0.775 mmol

mol(CH3NH2) = 4.5 mmol

0.775 mmol of both will react

excess CH3NH2 remaining = 3.725 mmol

Volume of Solution = 5 + 25 = 30 mL

[CH3NH2] = 3.725 mmol/30 mL = 0.1242 M

[CH3NH3+] = 0.775 mmol/30 mL = 0.0258 M

They form basic buffer

base is CH3NH2

conjugate acid is CH3NH3+

Kb = 4.4*10^-4

pKb = - log (Kb)

= - log(4.4*10^-4)

= 3.357

use:

pOH = pKb + log {[conjugate acid]/[base]}

= 3.357+ log {2.583*10^-2/0.1242}

= 2.675

use:

PH = 14 - pOH

= 14 - 2.6747

= 11.3253

Answer: 11.33